### v3

# Towards numerical two-loop integrand reduction

Abstract
We present a method for the integrand-level reduction of two-loop helicity amplitudes in both $d=4-2\epsilon$ and $d=4$ dimensions. The amplitude is expressed in terms of a set of Feynman integrals and their coefficients that depend on the external kinematics. The analysis presented in this paper, in conjunction with the ongoing development of the computational framework HELAC-2LOOP, paves the road for the construction of an automated program for two-loop amplitude calculations for arbitrary scattering processes.
Keywords:
Scattering Amplitudes, QCD Phenomenology, NNLO Computations, Integrand reduction
1 Introduction
Several decades after its establishment, the Standard Model (SM) keeps offering a successful description of Strong and Electroweak interactions among fundamental particles. One of the most concrete and reliable ways of probing the SM is the comparison of its predictions to experimental data from high-energy particle collisions. The present energy frontier is represented by the Large Hadron Collider (LHC) with its four experiments, CMS, ATLAS, ALICE, and LHCb. The forthcoming High Luminosity upgrade of the LHC (HL-LHC) will significantly boost the statistical accuracy of the experimental data that are being collected, allowing experiments to reach percent-level precision in many cases Dainese:2019rgk ; Caola:2022ayt ; CMS:2025hfp . The accurate interpretation of the data from the LHC, as well as from possible future colliders FCC:2025lpp , demands theoretical predictions of comparable precision.
Theoretical predictions are formulated from first principles within the framework of perturbative Quantum Field Theory, where scattering cross sections and other observables are computed as a power series expansion in the coupling constants of the SM. The first term in this expansion corresponds to the leading-order (LO) contribution, followed by subsequent corrections at next-to-leading order (NLO), next-to-next-to-leading order (NNLO), and so on. The calculation of NNLO corrections (and occasionally even N3LO) is necessary in order to meet the current (and near future) experimental precision. The bottleneck for most NNLO calculations consists, at present, of the computation of double-virtual corrections, which rely upon two-loop dimensionally regularized scattering amplitudes. As the number of particles in a process increases, the complexity of the latter increases dramatically. The current frontier is the calculation of $2\to 3$ processes Huss:2025nlt , with several results in this direction being achieved in recent years Badger:2017jhb ; Badger:2018enw ; Abreu:2018jgq ; Abreu:2018zmy ; Abreu:2019odu ; Badger:2019djh ; Hartanto:2019uvl ; Chawdhry:2020for ; Kallweit:2020gcp ; Abreu:2021oya ; Badger:2021nhg ; Badger:2021imn ; Badger:2021ega ; Chawdhry:2021mkw ; Chawdhry:2021hkp ; Czakon:2021mjy ; Abreu:2021asb ; Badger:2022ncb ; Badger:2023mgf ; Abreu:2023bdp ; Agarwal:2023suw ; DeLaurentis:2023nss ; DeLaurentis:2023izi ; Badger:2024sqv ; Badger:2024mir ; Mazzitelli:2024ura ; Badger:2024dxo ; Agarwal:2024jyq ; DeLaurentis:2025dxw as the product of considerable effort from the theoretical community.
The calculation of two-loop amplitudes can be regarded as a three-step procedure. The initial step involves the construction of the amplitude through the use of standard Feynman graph generation Hahn:2000kx ; Nogueira:1991ex or recursive approaches Pozzorini:2022ohr ; Canko:2023lvh . The second step includes the reduction of the amplitude into a set of independent Feynman integrals, commonly known as master integrals. For this to be achieved, the constructed amplitude should be projected into a set of scalar Feynman integrals, which can be done using tensor Peraro:2019cjj ; Peraro:2020sfm ; Anastasiou:2023koq ; Goode:2024mci ; Goode:2024cfy , or integrand Kosower:2011ty ; Mastrolia:2011pr ; Zhang:2012ce ; Badger:2012dp ; Mastrolia:2012an ; Mastrolia:2012wf ; Kleiss:2012yv ; Mastrolia:2013kca ; Badger:2013gxa ; Ita:2015tya ; Mastrolia:2016dhn ; Peraro:2019okx ; Abreu_2021_C reduction methods. The generated scalar integrals from these methods need to be furthermore reduced, at the integral level, into master integrals using integration-by-parts (IBP) identities Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2001dd , satisfied by the Feynman integrals within dimensional regularization ($d=4-2\epsilon$). The latter field has benefited from the application of finite-field techniques vonManteuffel:2014ixa ; Peraro:2016wsq for solving IBP identities, and the recent advancements in the generation of optimized IBP systems, utilizing syzygy equations Larsen:2015ped ; Wu:2023upw and the block-triangular form Liu:2018dmc ; Guan:2024byi . The last step for the computation of the scattering amplitude is the evaluation of the master integrals resulting from the reduction, which can be attained by methods such as sector decomposition Binoth:1999sp ; Heinrich:2008si ; Heinrich:2023til , differential equations Barucchi:1973zm ; Kotikov:1990kg ; Kotikov:1991hm ; Gehrmann:1999as ; Henn:2013pwa ; Moriello:2019yhu , auxiliary mass flow Liu:2017jxz ; Liu:2022chg , and dimension-changing transformation Huang:2024qan ; Huang:2024nij .
The present work focuses on the second step of the aforementioned procedure, and more specifically on the integrand reduction of two-loop scattering amplitudes into scalar integrals. In the one-loop case, the integrand reduction OPP method Ossola:2006us , demonstrated a systematic algebraic approach to the computation of amplitudes, based upon solving cut equations, meaning equations resulting from setting some (or all) of the inverse propagators to zero, following a top-down approach. Integrand reduction has been implemented in various numerical algorithms Ossola:2007ax ; Ellis:2007br ; Ellis:2008ir ; Berger:2008sj ; vanHameren:2009dr ; vanHameren:2009vq ; Mastrolia:2010nb ; Badger:2010nx ; Bevilacqua:2011xh ; Peraro:2014cba ; Alwall:2014hca ; GoSam:2014iqq ; Buccioni:2019sur and has boosted automation of NLO computations. Our goal is to extend the OPP method to two-loop calculations Kosower:2011ty ; Mastrolia:2011pr ; Zhang:2012ce ; Badger:2012dp ; Mastrolia:2012an ; Mastrolia:2012wf ; Kleiss:2012yv ; Mastrolia:2013kca ; Badger:2013gxa ; Ita:2015tya ; Mastrolia:2016dhn . We investigate two-loop amplitude reduction at the integrand level, in $d=4-2\epsilon$ and $d=4$ dimensions, by solving cut equations in a top-down approach Kosower:2011ty ; Mastrolia:2011pr ; Zhang:2012ce ; Badger:2012dp ; Badger:2013gxa ; Ita:2015tya ; Mastrolia:2016dhn ; Abreu_2021_C in several two-loop examples. In general, to facilitate computations, Feynman integrals that share the same topological and kinematical structure can be grouped into integral families. A novel element of our study is the extension of the cut equations to include all inverse propagators characterizing a provided integral family and not just the ones of the Feynman graph under consideration. This approach yields a larger number of systems of equations to be solved, albeit smaller in size, and has the advantage of allowing the reduction of all the Feynman graphs that belong to the same family altogether. Furthermore, we study an alternative method for integrand-level reduction that does not rely on cut equations, but rather on sampling the integrand’s numerator over random values of loop momenta.
This paper is structured as follows. In section 2, the notation employed is introduced, and the general ideas and concepts of the methods studied are explained. In section 3, the application of these methods to several examples with different kinematics is presented, ranging from four- to six-particle processes. In section 4, we conclude by giving a summary of our study and its prospects for the future.
2 Description of the method
2.1 Basic notation and conventions
A generic unrenormalized $L$-loop scattering amplitude can be written schematically as
${\cal A}^{(L)}(\{p\})=\sum_{g=1}^{G}\,\left(\int\prod_{i=1}^{L}\,[dk_{i}]\,\frac{{\cal N}(k_{1},\dots,k_{L},\{p\})}{D_{1}\,D_{2}\,\cdots D_{n}}\,\right)_{g}\,$, (1)
where $k_{1},\dots,k_{L}$ are loop momenta, $dk_{i}]$ denotes the integration measure associated to the loop momentum $k_{i}$, and $\{p\}$ represents the set of external kinematics. The amplitude consists generally of $G$ partial contributions, individually denoted by $g$. Depending on the considered approach, the latter may be individual Feynman diagrams or consistently defined sub-amplitudes. In the following, we will refer to them as loop topologies. In what follows, we focus on the genuine two-loop contributions to the amplitude (so-called Theta topologies), namely those that cannot be trivially reduced to the product of one-loop ones (Infinity and Dumbbell topologies) [^1] For the terminology, see Figure 1 of Ref. Bevilacqua:2024fec .. Each loop topology is characterized by a numerator $\cal N$}, which is a function of external momenta and wavefunctions as well as loop momenta, and by a product of $n$ inverse propagators $D_{1},\dots,D_{n}$, appearing in its denominator [^2] Inverse propagators raised in higher integer powers must also be included to accommodate special contributions..
The first step towards constructing a reduction method for two-loop calculations is to parametrize the numerator $\mathcal{N}$ in terms of the inverse propagators, moving along the path set by Ref. Ossola:2006us . A few comments are in order at this point. The numerator consists of independent scalar products involving loop and external momenta and possibly transverse vectors, i.e., $k_{i}\cdot k_{j}$, $k_{i}\cdot p_{j},k_{i}\cdot\eta_{j}$. The number $N$ of independent scalar products among loop and external momenta is easily derived, depending on the considered regularization scheme. Denoting with $L$ the number of loops and with $E$ the number of external momenta, in the Conventional Dimensional Regularization (CDR) scheme
N=L(L+1)/2+L(E1),N=L(L+1)/2+L(E-1)\,, (2)
whereas in the t’Hooft-Veltman (tHV) scheme
$N=\begin{cases}L(L+1)/2+L(E-1)&\text{for}\,\,\,E<5\\ L(L+1)/2+4L&\text{for}\,\,\,E\geq 5\end{cases}.$ (3)
In one-loop calculations ($L=1$), the number of inverse propagators $n$ appearing in the denominator of Eq. (1) is always larger than or equal to $N$. Consequently, all scalar products appearing in the numerator $\cal N$}, besides the so-called spurious terms Ossola:2006us , can be expressed in terms of the inverse propagators $D_{i}$, meaning they are all reducible scalar products (RSP). Conversely, at higher loops ($L\geq 2$), part of the aforementioned scalar products are not expressible in terms of the $D_{i}$’s. The latter are commonly denoted irreducible scalar products (ISP).
Extending the OPP approach to two loops, one can express the numerator $\mathcal{N}$ of a generic two-loop integrand in the following form:
${\cal N}=P^{(n)}+\sum_{i=1}^{n}P_{i}^{(n-1)}D_{i}+\sum_{i=1}^{n-1}\sum_{j>i}^{n}P_{ij}^{(n-2)}D_{i}D_{j}+\ldots+P_{12\ldots n}^{(0)}D_{1}D_{2}\ldots D_{n}$ (4)
The equation holds, as in the case of one loop Ossola:2006us , whenever the numerator of the amplitude has a rank, in $k_{1}$ and $k_{2}$, that does not exceed the number of corresponding propagators. In fact, as we will see later, the expansion may end long before its final contribution, depending on the rank of the numerator in question (see also Eq. (22) and section 3). Plugging this expression in Eq. (1) results in a splitting of the integrand into terms with a different number of denominators. The quantities $P^{(n)}$, $P_{i}^{(n-1)}$, $P_{ij}^{(n-2)},\dots$ appearing in Eq. (4) are polynomials of the form
$P=\sum_{l=1}^{M}\,b_{l}\,m_{l},\,$ (5)
where $b_{i}=b_{i}(\{p\})$ are coefficients which depend on the external kinematics, and the monomials $m_{i}$ are built upon the ISP characteristic of each term, and possibly of scalar products of the form $k_{i}\cdot\eta_{j}$.
Focusing on the two-loop case, our goal is to decompose the amplitude, ${\cal A}^{(2)}$, in terms of a set of Feynman integrals $F_{i}$ and coefficients $c_{i}$ that depend only on the external kinematics,
${\cal A}^{(2)}=\sum_{i}c_{i}(\{p\})\,F_{i}\,$, (6)
where the $F_{i}$’s take the form
$F_{i}\equiv F(a_{1},\ldots,a_{N})=\int\prod_{i=1}^{2}[dk_{i}]\frac{1}{D_{1}^{a_{1}}\ldots D_{N}^{a_{N}}},\;\;\;\;a_{i}\in\mathbb{Z}\,.$ (7)
This can be achieved by appropriately expressing the monomials $m_{l}$ in Eq. (5) in terms of the inverse propagators $D_{i}$ Sotnikov:2019onv . Eq. (6) is at the core of numerical methods for two-loop computations. Provided that the integrals $F_{i}$ are known (or a procedure to reduce the latter to a subset of master integrals is established), determining the coefficients $c_{i}$ at integrand level helps to address the problem of two-loop computations in a general, process-independent way. This idea expands upon well-established methods developed for one-loop calculations, such as OPP reduction Ossola:2006us .
In the following subsections, we will introduce two different methods that can be used to determine the polynomials $P$ appearing in Eq. (4) and to fit their coefficients.
2.2 Linear fit and fit by cut approach
We first address the question of determining the polynomials $P$ and fitting their coefficients in Eq. (4), assuming that the numerator is known analytically. This allows us to construct the polynomials $P$ directly from the symbolic expressions appearing in the numerator. Moving along the same path of one-loop reduction, the fitting is carried out by iteratively applying cut equations, namely setting denominators to zero. Starting with the so-called maximal cut equation,
$D_{1}=D_{2}=\ldots=D_{n}=0\,$, (8)
we can identify the coefficients of $P^{(n)}$,
$P^{(n)}={\cal N}|_{D_{1}=D_{2}=\ldots=D_{n}=0}\,.$ (9)
Then we can iteratively fit the rest of the polynomials by appropriately subtracting the terms computed in the previous step. For instance, the first next-to-maximal contribution reads
$P^{(n-1)}_{1}=\left.\left(\frac{{\cal N}-{\cal N}|_{D_{1}=D_{2}=\ldots=D_{n}=0}}{D_{1}}\right)\right\rvert_{D_{2}=\ldots=D_{n}=0}\,$, (10)
and so on. We note that cut equations, Eq. (8), translate into a system of linear relations among scalar products $k_{i}\cdot k_{j}$ and $k_{i}\cdot p_{j}$, where $k_{i}$ and $p_{j}$ denote generically loop and external momenta, respectively. This allows us to straightforwardly solve Eq. (9), Eq. (10), and the rest of the equations resulting from all the sub-maximal cuts, by substitution rules, for any process. We call this procedure of solving Eq. (4), a linear fit. We will show in the following sections how this applies to the case of 4-, 5-, and 6-particle scattering amplitudes.
In certain instances, the analytical calculation of the numerator $\cal N$} can prove to be a highly challenging task. In such cases, resorting to a numerical computation, facilitated by dedicated software packages such as HELAC-2LOOP Canko:2023lvh , can be a feasible alternative. In this case, there are two issues to be addressed:
  1. 1.
    the solutions of cut equations must be expressed in a form suitable for numerical evaluation of numerators;
  2. 2.
    the polynomials $P$ appearing in Eq. (4) must be constructed without a priori analytical knowledge of the numerator.
To address the first issue, we need a suitable representation of the loop momenta. Given two arbitrary massless momenta $l_{1}^{\mu},l_{2}^{\mu}$, let us define
$\begin{gathered}l_{3}^{\mu}=\bar{u}_{-}(l_{1})\gamma^{\mu}u_{-}(l_{2})\,,\\ l_{4}^{\mu}=\bar{u}_{-}(l_{2})\gamma^{\mu}u_{-}(l_{1})\,.\end{gathered}$ (11)
The set $\{l_{1}^{\mu},l_{2}^{\mu},l_{3}^{\mu},l_{4}^{\mu}\}$ forms a basis in $d=4$ dimensions. This allows us to express the loop momenta $k_{1},k_{2}$ as follows,
$\begin{gathered}k_{1}=x_{1}\,l^{(1)}_{1}+x_{2}\,l^{(1)}_{2}+x_{3}\,l^{(1)}_{3}+x_{4}\,l^{(1)}_{4}\\ k_{2}=y_{1}\,l^{(2)}_{1}+y_{2}\,l^{(2)}_{2}+y_{3}\,l^{(2)}_{3}+y_{4}\,l^{(2)}_{4}\end{gathered}$ (12)
where the coefficients $x_{i}$ and $y_{i}$ are expressible in terms of scalar products of the form $k_{i}\cdot p_{j}$. The latter coefficients characterize the loop momenta in $d=4$ dimensions. Complemented by $\mu_{11}$, $\mu_{12}$ and $\mu_{22}$ [^3] In $d=4-2\epsilon$, $\mu_{ij}\equiv k^{(\epsilon)}_{i}\cdot k^{(\epsilon)}_{j}$, with $k^{(\epsilon)}$ the components of the loop momenta beyond $d=4$ dimensions., they form a set of eleven variables which characterizes completely the loop momenta in $d=4-2\epsilon$ dimensions: $\vec{X}=\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},y_{3},y_{4},\mu_{11},\mu_{12},\mu_{22}\}$. We solve cut equations [^4] For alternative approaches, see Refs. Kosower:2011ty ; Mastrolia:2011pr ; Zhang:2012ce ; Badger:2012dp ; Badger:2013gxa ; Ita:2015tya . in terms of these variables, as we will see in section˜3. In $d=4-2\epsilon$, the solution to the cut equations is unique in terms of the ISP, whereas in $d=4$ we usually have disjoint branches Zhang:2012ce , see section˜3.1.1 for an explicit example.
The parametrization of the polynomials $P$ in terms of the ISP is obtained through the program BasisDet Zhang:2012ce . The latter provides a set of monomials which take the form $\prod_{i}x_{i}^{r_{i}}$, where $x_{i}$ denote ISP and $r_{i}$ is an integer ranging from zero to some upper value calculated from the maximal tensor rank of the polynomial $P$ with respect to $k_{1}$, $k_{2}$ and $k_{1},k_{2}$ combined [^5] Corresponding to RenormalizationCondition in ref. Zhang:2012ce .. For example, a term such as $(k_{1}\cdot p_{1})^{2}(k_{2}\cdot p_{3})^{3}$ has a tensor rank 2 for $k_{1}$, 3 for $k_{2}$ and 5 for $k_{1},k_{2}$ combined.
At one loop, $P$ consists of terms depending solely on the external kinematics and the so-called spurious terms, which are specific to each cut. The spurious terms, although necessary for the reduction at the integrand level, do not contribute to the final result as they integrate to zero. The final result is determined by the coefficients that depend only on the external kinematics and multiply the appropriately chosen basis of integrals. At two loops, the existence of spurious terms that integrate to zero is less straightforward: there are certainly spurious terms compiled by the loop momenta and the transverse directions over the external momenta, whenever present. Nevertheless, the simple one-loop picture is spoiled by the fact that the integrals in Eq. (7) obey a set of IBP identities, resulting in a set of master integrals, which are then evaluated using different techniques. In this paper, we do not address the issue of IBP reduction: we assume that all the integrals resulting from the integrand reduction are, or will be, available numerically, as happens for the one-loop case.
[…]
${\cal N}\equiv{\cal N}(\vec{X},d)={\cal N}_{0}+\sum_{i\geq 1}\epsilon^{i}{\cal N}^{(i)}_{\epsilon}\,$, (13)
by expanding in powers of $\epsilon\equiv(4-d)/2$, where both ${\cal N}_{0}\equiv{\cal N}(\vec{X},d)|_{d=4}$ and ${\cal N}^{(i)}_{\epsilon}\equiv{\cal N}^{(i)}_{\epsilon}(\vec{X})$, are accessible numerically and depending on $\mu_{ij}$ through $\vec{X}$. These terms are used to calculate the $M\times 1$ matrices, ${\cal B}^{(0)}_{j}={\cal N}_{0}(\vec{X}_{j})$, ${\cal B}^{(i)}_{j}={\cal N}^{(i)}_{\epsilon}(\vec{X}_{j})$. Then the given polynomial $P$ is written explicitly as
$P=\sum_{i=1}^{M}\left(c_{i}^{(0)}+\sum\epsilon^{j}\;c_{i}^{({j})}\right)m_{i}$ (14)
where
${\vec{c}\;}^{(0)}={\cal M}^{-1}{\cal B}^{(0)}\;\;\;\;\;{\vec{c}\;}^{({i})}={\cal M}^{-1}{\cal B}^{(i)}\;\;\;\;\;$ (15)
After the whole iterative procedure is completed, the so-called $N=N$ test is performed. The latter consists of checking the validity of Eq. (4) for arbitrary assignment of numerical values for all free parameters of the loop kinematics, $\vec{X}$, not restricted by cut equations. We have checked that the $N=N$ test is fulfilled when the polynomials $P$ are constructed directly from the analytic expression of the numerator, as well as when using BasisDet to construct the ansatz for the polynomials.
[…]
${\cal M}\,{\vec{c}\;}^{(d=4)}={\cal B}_{0}^{(4)}$ (16)
can still be solved with standard Linear Algebra algorithms such as QR decomposition, as long as the rank of the matrix is full, namely $rank({\cal M})=M$. As we will see later, this is true in most cases, but solutions can still be obtained in cases where the matrix is rank-deficient, $rank({\cal M})<M$. Notice that in $d=4-2\epsilon$ case, when the information on the dependence of the numerator on $d$ and $\mu_{ij}$ is available, the reduction of the amplitude is complete, whereas in four dimensions, where this information is not available, the so-called rational terms need to be calculated in addition Ossola:2008xq ; Badger:2008cm ; Pozzorini:2020hkx ; Lang:2020nnl ; Lang:2021hnw .
2.3 Global fit approach
[…]
2.4 Projecting over a full family
From the perspective of the one-loop OPP approach, Eq. (4) addresses the reduction of the numerator in Eq. (1) in terms of the $n$ inverse propagators $D_{i}$ appearing in it. As we already pointed out, the one-loop case is special in the sense that the number of independent scalar products $N$ and the number of inverse propagators $n$ obey the relation $N\leq n$. Thus, all scalar products can be expressed in terms of the $D_{i}$’s, which appear in the denominator of the loop integrand. Starting from two loops, $N>n$, and thus one is left with a set of ISP that cannot be expressed as above. However, one can define an enlarged set of inverse propagators such that all scalar products are expressible as combinations of the latter. This enlarged set of inverse propagators is named family. We can consider projecting the numerator over the full family of inverse propagators:
${\cal N}=P^{(N)}+\sum_{i=1}^{N}P_{i}^{(N-1)}D_{i}+\sum_{i=1}^{N-1}\sum_{j>i}^{N}P_{ij}^{(N-2)}D_{i}D_{j}+\ldots+P_{12\ldots N}^{(0)}D_{1}D_{2}\ldots D_{N}$ (17)
For example, cutting over 7 propagators in a two-loop four-particle topology,
$\begin{gathered}{\cal N}_{4p}=P_{7c}(v_{1},v_{2},u_{1},u_{2})+\sum_{i=1}^{7}P_{6c}^{(i)}(v_{1},v_{2},u^{i}_{1},u^{i}_{2},u^{i}_{3})D_{i}\\ +\sum_{i<j}^{7}P_{5c}^{(i,j)}(v_{1},v_{2},u^{i,j}_{1},u^{i,j}_{2},u^{i,j}_{3},u^{i,j}_{4})D_{i}D_{j}\\ +\sum_{i<j<k}^{7}P_{4c}^{(i,j,k)}(v_{1},v_{2},u^{i,j,k}_{1},\ldots,u^{i,j,k}_{5})D_{i}D_{j}D_{k}\\ +\sum_{i<j<k<l}^{7}P_{3c}^{(i,j,k,l)}(v_{1},v_{2},u^{i,j,k,l}_{1},\ldots,u^{i,j,k,l}_{6})D_{i}D_{j}D_{k}D_{l}\\ +\sum_{i<j<k<l<m}^{7}P_{2c}^{(i,j,k,l,m)}(v_{1},v_{2},u^{i,j,k,l,m}_{1},\ldots,u^{i,j,k,l,m}_{7})D_{i}D_{j}D_{k}D_{l}D_{m}+\ldots\end{gathered}$ (18)
with $v_{i}\equiv k_{i}\cdot\eta$, where $\eta$ is the vector transverse to $p_{1},p_{2},p_{3}$ necessary to complete the 4-dimensional physical space, while $u^{i}_{I}$, $u^{i,j}_{I}$, $u^{i,j,k}_{I}$, $u^{i,j,k,l}_{I},\,\dots$, stand for ISP in the corresponding cut, will generically lead to polynomials $P$ depending on more variables rather than cutting over all 9 propagators,
$\begin{gathered}{\cal N}_{4p}=P_{9c}(v_{1},v_{2})+\sum_{i=1}^{9}P_{8c}^{(i)}(v_{1},v_{2},u_{i})D_{i}+\sum_{i<j}^{9}P_{7c}^{(i,j)}(v_{1},v_{2},u_{i},u_{j})D_{i}D_{j}\\ +\sum_{i<j<k}^{9}P_{6c}^{(i,j,k)}(v_{1},v_{2},u_{i},u_{j},u_{k})D_{i}D_{j}D_{k}\\ +\sum_{i<j<k<l}^{9}P_{5c}^{(i,j,k,l)}(v_{1},v_{2},u_{i},u_{j},u_{k},u_{l})D_{i}D_{j}D_{k}D_{l}\\ +\sum_{i<j<k<l<m}^{9}P_{4c}^{(i,j,k,l,m)}(v_{1},v_{2},u_{i},u_{j},u_{k},u_{l},u_{m})D_{i}D_{j}D_{k}D_{l}D_{m}+\ldots\end{gathered}$ (19)
where now $u_{i}$ stand for ISP in the corresponding cut (see explicit expressions in section 3).
The same is true for a five-particle amplitude, where now the physical phase-space is fully described by the four external independent momenta, cutting over 8 propagators results in
$\begin{gathered}{\cal N}_{5}=P_{8c}(u_{1},u_{2},u_{3})+\sum_{i=1}^{8}P_{7c}^{(i)}(u^{i}_{1},\ldots,u^{i}_{4})D_{i}\\ +\sum_{i<j}^{8}P_{6c}^{(i,j)}(u^{i,j}_{1},\ldots,u^{i,j}_{5})D_{i}D_{j}\\ +\sum_{i<j<k}^{8}P_{5c}^{(i,j,k)}(u^{i,j,k}_{1},\ldots,u^{i,j,k}_{6})D_{i}D_{j}D_{k}\\ +\sum_{i<j<k<l}^{8}P_{4c}^{(i,j,k,l)}(u^{i,j,k,l}_{1},\ldots,u^{i,j,k,l}_{7})D_{i}D_{j}D_{k}D_{l}\\ +\sum_{i<j<k<l<m}^{8}P_{3c}^{(i,j,k,l,m)}(u^{i,j,k,l,m}_{1},\ldots,u^{i,j,k,l,m}_{8})D_{i}D_{j}D_{k}D_{l}D_{m}+\ldots\end{gathered}$ (20)
while cutting over all 11 propagators, which defines the 5-point family, results in
$\begin{gathered}{\cal N}_{5}=P_{11c}+\sum_{i=1}^{11}P_{10c}^{(i)}(u_{i})D_{i}+\sum_{i<j}^{11}P_{9c}^{(i,j)}(u_{i},u_{j})D_{i}D_{j}\\ +\sum_{i<j<k}^{11}P_{8c}^{(i,j,k)}(u_{i},u_{j},u_{k})D_{i}D_{j}D_{k}\\ +\sum_{i<j<k<l}^{11}P_{7c}^{(i,j,k,l)}(u_{i},u_{j},u_{k},u_{l})D_{i}D_{j}D_{k}D_{l}\\ +\sum_{i<j<k<l<m}^{11}P_{6c}^{(i,j,k,l,m)}(u_{i},u_{j},u_{k},u_{l},u_{m})D_{i}D_{j}D_{k}D_{l}D_{m}+\ldots\end{gathered}$ (21)
Notice that in this case $P_{11c}$ depends only on the external kinematics, since the maximal cut determines all degrees of freedom of the loop momenta. We observe that projecting over the full family has the effect of simplifying the structure of the individual polynomials $P$, since they depend on a reduced number of ISP. Moreover, all the ISP $u_{i}$ in Eq. (21), can be expressed in terms of the inverse propagators of the family, leading directly to a variation of the master equation,
${\cal N}_{5}=c^{(0)}+\sum_{i=1}^{11}\sum_{a}c^{(i,a)}D^{a}_{i}+\sum_{i<j}^{11}\sum_{a,b}c^{(i,j,a,b)}D_{i}^{a}D_{j}^{b}+\ldots$ (22)
where, first the coefficients $c$ depend only on the external kinematics and second the inverse Feynman propagators, $D_{i}$, enter not linearly as before, but with powers determined by the fit-by-cut procedure. The amplitude is then given by,
${\cal A}^{(2)}_{5}=c^{(0)}F[a_{1},\ldots,a_{11}]+\sum_{i=1}^{11}\sum_{a}c^{(i,a)}F[a_{1},\ldots,a_{i}-a,\ldots,a_{11}]+\ldots$ (23)
with $a_{1}=\ldots=a_{8}=1$ and $a_{9}=a_{10}=a_{11}=0$ and $F$ given as in Eq.(7). Eqs. (20-23) are straightforwardly extensible to any $n$-particle amplitude with $n>5$, see section 3.3.1 for details in case $n=6$.
3 Application of the method
In this section, the application of the method described in section˜2 is demonstrated through several examples. The considered numerators are representative of a variety of scattering processes and consist of different kinematic dependencies, ranging from four-point to six-point kinematics. For the generation of the analytic expressions required, we used the Mathematica packages FeynArts Hahn:2000kx and FeynCalc Shtabovenko:2023idz , except for the 6-particle case, section 3.3.1, which has been generated by FORM Ruijl:2017dtg .
3.1 Four-Point Kinematics
3.1.1 Double-box topologies
Figure 1: Feynman graphs contributing to the double-box numerator under study. There are seven contributions, considering gluons and (anti-)ghosts running within the loops. Curly lines denote gluons, while dotted lines denote ghosts (arrow aligned with the momentum flow) and anti-ghosts (arrow in the opposite direction of the momentum flow).
Here we focus on the numerator of the double-box topology constructed by the seven Feynman graphs depicted in Fig. 1. This numerator contributes to the scattering amplitude of the process $gg\to gg$. The inverse propagators describing the family for this topology can be chosen as
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{3}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+k_{2}\right){}^{2},\quad D_{5}=k_{2}^{2},\\ D_{6}=\left(k_{2}-p_{123}\right){}^{2},\quad D_{7}=\left(k_{2}-p_{12}\right){}^{2},\quad D_{8}=\left(k_{2}-p_{1}\right){}^{2},\quad D_{9}=\left(k_{1}+p_{123}\right){}^{2}\end{gathered}\,$ (24)
Above and henceforth, the shorthand notation $p_{i\ldots j}=p_{i}+\ldots+p_{j}$ will be used to denote the sum of the incoming external on-shell momenta, and $s=(p_{1}+p_{2})^{2}$ and $t=(p_{2}+p_{3})^{2}$ the standard Mandelstam variables. For all analyses presented below we have used the following helicity assignment: $\lambda_{1}=+1,\lambda_{2}=+1,\lambda_{3}=-1,\lambda_{4}=-1$, for the helicities of the external gluons, all assumed to be incoming.
  1. 1.
    Linear fit
    We seek to solve Eq. (18) using cut equations and the known analytic expression of the numerator.
    The maximal cut equations
    $D_{1}=D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=0$ (25)
    result in determining seven invariants
    $\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot k_{2}\to 0,\\ k_{2}\cdot p_{2}\to\frac{s}{2}-k_{2}\cdot p_{1},\quad k_{2}\cdot p_{3}\to-\frac{s}{2}.\end{gathered}$ (26)
    By applying the above relation on both sides of Eq. (18), we can fully determine the polynomial $P_{7}$. The latter consists of 70 coefficients over the ISP monomials $\{k_{1}\cdot p_{3},\;k_{1}\cdot\eta,\;k_{2}\cdot p_{1},\;k_{2}\cdot\eta\}$.
    Subtracting $P_{7}$ in Eq. (18), we can now determine the polynomials of kind $P_{6}$ in the same way. There are seven six-cuts and therefore seven $P_{6}$ polynomials to determine. As an example, the first six-cut,
    $D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=0$ (27)
    leads to
[…]
    where there are now 5 ISP: $\{k_{1}\cdot p_{2},\;k_{1}\cdot p_{3},\;k_{1}\cdot\eta,\;k_{2}\cdot p_{1},\;k_{2}\cdot\eta\}$. The polynomial $P^{(6)}_{1}$ consists of 111 coefficients. This process is iterated until the level of a two-cut, after which all resulting polynomials vanish. The data for all cuts are summarized in Tab. 1. The analytic solution for the polynomials satisfies explicitly Eq. (18).
    Level Number of cuts Number of coefficients Scaling
    7 1 70 4,4,4
    6 7 695 3,3,4
    5 21 1430 3,3,3
    4 35 1017 2,2,2
    3 35 225 1,1,1
    2 21 9 0,0,0
    Table 1: Double-box linear fit information beginning with 7-cut. The numbers in the last column refer to the maximum powers of $k_{1}$, $k_{2}$, and $k_{1},k_{2}$ combined, as described in the text.
    We now seek to solve Eq. (19), namely projecting over all 9 propagators in the double-box family. This has the advantage of building a reduction procedure that covers all 4-particle planar diagrams, including the double-box, the penta-triangle and the hexa-bubble. The maximal cut equations read
    $D_{1}=D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=D_{8}=D_{9}=0$ (29)
    which leads to
    $\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\\ k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2}\end{gathered}$ (30)
    By applying the above relations on both sides of Eq. (19), we can fully determine the polynomial $P_{9}$. The latter consists of 13 coefficients over the ISP monomials $\{k_{1}\cdot\eta,\;k_{2}\cdot\eta\}$. Subtracting as before $P_{9}$ in Eq. (19), we can now determine the polynomials $P_{8}$ in the same way. There are nine 8-cuts and therefore nine $P_{8}$ polynomials to determine. As an example, the first 8-cut,
    $D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=D_{8}=D_{9}=0$ (31)
    leads to
[…]
    where there are now 3 ISP: $\{k_{1}\cdot p_{1},\;k_{1}\cdot\eta,\;k_{2}\cdot\eta\}$.
    Level Number of cuts Number of coefficients Scaling
    9 1 13 4,4,4
    8 9 227 4,4,4
    7 36 963 3,3,3
    6 84 1445 2,2,2
    5 126 780 1,1,1
    4 126 116 0,0,0
    Table 2: Double-box linear fit information beginning with 9-cut.
    The data for all cuts are summarized in Tab. 2. The analytic solutions for the polynomials satisfy explicitly Eq. (18). The total number of non-zero coefficients is slightly larger than previously, namely 3544 versus 3446. We have verified that after reducing by IBP identities the integrals appearing in Eqs. (18,19), the coefficients of the top-sector master integrals coincide with those obtained from Caravel Abreu_2021_C , for all helicity assignments  [^6] In Caravel the results are given for the color-stripped helicity amplitude, whereas in our case we have studied a subset of the contributions to the amplitude, as given in Fig. 1. Nevertheless, a comparison of the top-sector master integral coefficients is possible since all other Feynman graphs do not contribute to them.
  2. 2.
    Fit by cut in $d=4-2\epsilon$ dimensions
    Let us now assume that the numerator is available only using a numerical approach, as the one implemented in HELAC-2LOOP, including terms proportional to $\mu_{ij}$ and $\epsilon=(d-4)/2$. Then the realization of the solutions of the cut equations, Eq. (26) in a numerical setup is based on the determination of the four-dimensional part of the loop momenta following Eq. (12). In fact the solution for any cut has a unique analytic form in terms of ISP. The 7 cut, Eq. (26), reads as follows:
    $\begin{gathered}k_{1}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{1}{2}\left(s-2k_{2}\cdot p_{1}\right),\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2},\\ \mu_{11}\to-\frac{\frac{4s\left(k_{1}\cdot p_{3}\right){}^{2}}{s+t}-4sk_{1}\cdot p_{3}+4t\left(k_{1}\cdot\eta\right){}^{2}+s(s+t)}{4t},\\ \mu_{12}\to\frac{\frac{k_{1}\cdot p_{3}\left(4(s+2t)k_{2}\cdot p_{1}-2st\right)}{s+t}+t\left(s-4k_{1}\cdot\eta\,k_{2}\cdot\eta\right)-2sk_{2}\cdot p_{1}}{4t},\\ \mu_{22}\to-\frac{4stk_{2}\cdot p_{1}+4s\left(k_{2}\cdot p_{1}\right){}^{2}+t\left(4(s+t)\left(k_{2}\cdot\eta\right){}^{2}+st\right)}{4t(s+t)}\end{gathered}$ (33)
    For comparison, the cut conditions for 9 propagators, Eq. (30), are
    $\begin{gathered}k_{2}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s}{2},\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\\ k_{2}\cdot p_{3}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\quad\mu_{11}\to-\left(k_{1}\cdot\eta\right){}^{2}-\frac{st}{4(s+t)},\\ \quad\mu_{12}\to\frac{st}{4(s+t)}-k_{1}\cdot\eta k_{2}\cdot\eta,\quad\mu_{22}\to-\left(k_{2}\cdot\eta\right){}^{2}-\frac{st}{4(s+t)}\end{gathered}$ (34)
    In terms of the basis introduced in section 2.2, i.e. $\{x_{1},\dots,x_{4},y_{1},\dots,y_{4},\mu_{11},\mu_{12},\mu_{22}\}$, the solution takes the form
    $\begin{gathered}x_{1}\to-1,\quad x_{2}\to 0,\quad y_{1}\to 0,\quad y_{2}\to-1,\quad\mu_{11}\to 4sx_{3}x_{4},\quad\mu_{22}\to 4sy_{3}y_{4},\\ \mu_{12}\to(x_{3}+x_{4}-y_{3}-y_{4})r-2(s+t)(x_{4}y_{4}+x_{3}y_{3})-2t\left(x_{3}y_{4}+x_{4}y_{3}\right)-t/2\end{gathered}$ (35)
    with $r=\sqrt{-t(s+t)}$. In a numerical setup, for instance, calculating on the kinematic point s=1,t=1/5s=1,\,t=-1/5, the 7-cut is represented by
    $\begin{gathered}x_{1}\to-1,\quad x_{2}\to 0,\quad y_{1}\to 0,\quad y_{2}\to-1,\quad\mu_{11}\to 4x_{3}x_{4},\quad\mu_{22}\to 4y_{3}y_{4},\\ \mu_{12}\to\left(4x_{3}\left(-4y_{3}+y_{4}+1\right)+4x_{4}\left(y_{3}-4y_{4}+1\right)-4y_{3}-4y_{4}+1\right)/10\end{gathered}$ (36)
    We can now determine the coefficients of the polynomial $P_{7}$, which in this case has 70 coefficients, by calculating the numerator $\cal N$}, and the monomials $m_{i}$ of the basis, using 70 random assignments of the undetermined variables, $x_{3},x_{4},y_{3},y_{4}$, i.e. ${\cal N}_{i}$ and $M_{ij}$ respectively, and solving the corresponding matrix equation
    $\sum_{j=1}^{70}M_{ij}\,c_{j}={\cal N}_{i},\;\;\;i=1,\ldots,70$ (37)
    for the unknown coefficients $c$. The full-rank matrix $M$ is straightforwardly invertible, and the solution checked agrees to the numerical precision used against the analytic result. We have confirmed that this way we can calculate all the coefficients of Tab. 1 numerically. The same is true for the case of projecting over the 9 propagators, see Tab. 2.
  3. 3.
    Fit by cut in $d=4$ dimensions
    Level RenormalizationCondition
    7 {{{1,0},4},{{0,1},4},{{1,1},4}}
    6 {{{1,0},4},{{0,1},4},{{1,1},4}}
    5 {{{1,0},3},{{0,1},3},{{1,1},3}}
    4 {{{1,0},2},{{0,1},2},{{1,1},2}}
    3 {{{1,0},1},{{0,1},1},{{1,1},1}}
    2 {{{1,0},0},{{0,1},0},{{1,1},0}}
    Table 3: Input used for BasisDet, denoting by $\{\{1,0\},n_{1}\}$ the maximal power $n_{1}$ in $k_{1}$, by $\{\{0,1\},n_{2}\}$ the maximal power $n_{2}$ in $k_{2}$ and by $\{\{1,1\},n_{12}\}$ the maximal power $n_{12}$ in $k_{1}$, $k_{2}$ combined.
    In this case, the monomials that make up the polynomials of the analytic solution described either by the linear fit or the fit by cut in $d=4-2\epsilon$, are not independent, due to the existence of Gram-determinant identities which hold in $d=4$ dimensions. This is why for $d=4$ we use BasisDet as a means to parametrize the polynomials, therefore no longer relying on analytical knowledge of the numerator in order to create the anzatz. By construction, this requires us to specify the highest powers of the loop momenta to the BasisDet program. Following the data of Tab. 1, the so-called RenormalizationCondition in BasisDet is fixed as shown in Tab. 3. Contrary to the $d=4-2\epsilon$ case discussed previously, in $d=4$ the cut equations do not admit a unique solution in terms of ISP. Using the same values for $s,t$ as above, the 7-cut solution is split into 6 branches,
    $\begin{gathered}\left\{x_{1}\to-1,x_{2}\to 0,x_{3}\to 0,x_{4}\to\frac{4y_{3}-1}{4\left(y_{3}+1\right)},y_{1}\to 0,y_{2}\to-1,y_{4}\to 0,y_{3}\to z(1)\right\},\\ \left\{x_{1}\to-1,x_{2}\to 0,x_{3}\to\frac{4y_{4}-1}{4\left(y_{4}+1\right)},x_{4}\to 0,y_{1}\to 0,y_{2}\to-1,y_{3}\to 0,y_{4}\to z(1)\right\},\\ \left\{x_{1}\to-1,x_{2}\to 0,x_{3}\to 0,x_{4}\to-\frac{1}{4},y_{1}\to 0,y_{2}\to-1,y_{3}\to 0,y_{4}\to z(1)\right\},\\ \left\{x_{1}\to-1,x_{2}\to 0,x_{3}\to-\frac{1}{4},x_{4}\to 0,y_{1}\to 0,y_{2}\to-1,y_{4}\to 0,y_{3}\to z(1)\right\},\\ \left\{x_{1}\to-1,x_{2}\to 0,x_{3}\to 0,y_{1}\to 0,y_{2}\to-1,y_{3}\to 0,y_{4}\to\frac{1}{4},x_{4}\to z(1)\right\},\\ \left\{x_{1}\to-1,x_{2}\to 0,x_{4}\to 0,y_{1}\to 0,y_{2}\to-1,y_{3}\to\frac{1}{4},y_{4}\to 0,x_{3}\to z(1)\right\},\end{gathered}$
    each one parametrized by one undetermined variable $z(1)$. The polynomial $P_{7}$ is parametrized by only 28 coefficients, in $d=4$. By assigning 28 random values to $z(1)$ for each solution, we may form a matrix equation as follows:
    $\sum_{j=1}^{28}M_{ij}\,c_{j}={\cal N}_{i},\;\;\;i=1,\ldots,168$ (38)
    where the rows of the matrix $M$ consist of the values of the monomials involved in the basis and $\cal N$ are the corresponding values of the numerator. Although this is a rectangular system of equations, it admits a unique solution for the coefficients $c$. We have checked against the analytic solution and found agreement up to the numerical precision used. The same is true for all lower-level cuts.
  4. 4.
    Global fit in $d=4-2\epsilon$ dimensions
    As an alternative to determining the structure of Eq. (18), through cut equations, we have also studied the global fit, as described in the previous section.By expressing both sides of Eq. (18) in terms of the variables $V=\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},y_{3},y_{4},\mu_{11},$ $\mu_{12},\mu_{22}\}$, parameterizing the loop momenta, we can form a matrix equation over all undetermined parameters $c$ entering in the definition of the polynomials $P_{7},\ldots,P_{2}$ in Eq. (18). The number of parameters $c$, as well as the monomials involved, is 3446. Assigning random values to all 11 variables of $V$, we form the equation
    $\sum_{j=1}^{3446}M_{ij}\,c_{j}={\cal N}_{i},\;\;\;i=1,\ldots,3446$ (39)
    where, as before, $M$ is the matrix formed by the numerical values of the monomials and $\cal N$ the numerical values of the numerator. The matrix $M$ is of full rank and the solution has been determined both in Mathematica as well as using purely numerical tools, such as Eigen and LAPACK. In fact, the latter offers us the possibility to obtain our results in both double and quadruple precision. As a check of the validity of the solution, the $N=N$ test is successfully performed, in full precision, for arbitrary random assignment of the variables in $V$.
  5. 5.
    Global fit in $d=4$ dimensions
    A global fit in $d=4$ can be performed following the same line of reasoning as before. The total number of coefficients $c$ and monomials is 3018. The variables describing the loop momenta are $V=\left\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},y_{3},y_{4}\right\}$. Assigning random values to all 8 variables of $V$, we form the equation
    $\sum_{j=1}^{3018}M_{ij}c_{j}={\cal N}_{i},\;\;\;i=1,\ldots,3018$ (40)
    where as before $M$ is the matrix formed by the numerical values of the monomials and $\cal N$ the numerical values of the numerator in $d=4$ dimensions. The matrix $M$ is of full rank and the solution has been determined both in Mathematica as well as using purely numerical engines, such us Eigen and LAPACK. In fact the numerical engines offer us the possibility to obtain our results in both double and quadruple precision. As a check of the validity of the solution, the $N=N$ test is successfully performed, in full precision, for arbitrary random assignment of the variables in $V$.
3.1.2 Penta-triangle topologies
Figure 2: Feynman graphs contributing to the penta-triangle numerator under study. There are seven contributions, considering gluons and (anti-)ghosts running within the loops.
In this subsection, we study the numerator of the penta-triangle topology produced by the seven Feynman graphs depicted in Fig. 2. This numerator is part of the scattering amplitude of the process $gg\to gg$. The helicity assignment is as in the case of double-box. The propagators describing the family for this topology coincide with the ones of Eq. (24), but we choose a different ordering such that the auxiliary propagators appear in the end
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{2}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+p_{123}\right){}^{2},\quad D_{5}=\left(k_{1}+k_{2}\right){}^{2},\\ D_{6}=k_{2}^{2},\quad D_{7}=\left(k_{2}-p_{123}\right){}^{2},\quad D_{8}=\left(k_{2}-p_{12}\right){}^{2},\quad D_{9}=\left(k_{2}-p_{1}\right){}^{2}\end{gathered}$
  1. 1.
    Linear fit, fit by cut, and global fit in $d=4-2\epsilon$ dimensions
    The $d=4-2\epsilon$ analysis is producing the same qualitative results as in the double-box case. For instance, the 7-cut is given by
    $\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\\ k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{3}\to-\left(k_{2}\cdot p_{1}\right)-k_{2}\cdot p_{2}\end{gathered}$ (41)
    The data regarding the number of cuts, coefficients of polynomials, and respective powers are collected in Tab. 4.
    Level Number of cuts Number of coefficients Scaling
    7 1 53 4,4,4
    6 7 573 4,4,4
    5 21 1249 3,3,3
    4 35 929 2,2,2
    3 35 206 1,1,1
    2 21 8 0,0,0
    Table 4: Penta-triangle linear fit information beginning with 7-cut.
    The corresponding data when projecting over 9 propagators are shown in Tab. 5.
    Level Number of cuts Number of coefficients Scaling
    7 1 9 3,3,3
    6 9 573 4,3,4
    5 36 899 3,3,3
    4 84 1279 2,2,2
    3 126 652 1,1,1
    2 126 100 0,0,0
    Table 5: Penta-triangle linear fit information beginning with 9 cut.
  2. 2.
    Fit by cut and global fit $d=4$
    In $d=4$, as before, we have to rely on BasisDet in order to construct monomial bases. In this case, the numerator evaluated on all 7-cut solutions is zero. In the 6-cut sector, a new issue arises, namely, there are cuts for which the corresponding matrix equation, see Eq. (38), is rank deficient, meaning that the solution is described in terms of undetermined coefficients. These undetermined coefficients, however, do not impact the $N=N$ test. The same is true for the global fit, see Eq. (40), where again the equation is rank deficient, nevertheless, its solution correctly describes the numerator. We note that the penta-triangle is the only case studied in this paper where such a problem occurs.
[…]
3.1.3 Hexa-bubble topologies
Figure 3: Feynman graphs contributing to the hexa-bubble numerator under study. There are four contributions, considering gluons and (anti-)ghosts running within the loops.
We now consider the numerator of the hexa-bubble topology consisting of the four Feynman graphs depicted in Fig. 3. This numerator is part of the scattering amplitude of the process $gg\to gg$. The helicity assignment is as in the case of double-box. The inverse propagators describing this topology coincide with the ones of penta-triangle topology, but only the first six of those propagators appear in the denominator, with the first one being repeated and thus being raised to the second power. The analysis, projecting over 6 or 9 propagators, works for all cases, as in the double-box, in both $d=4-2\epsilon$ and $d=4$ dimensions. Of course, in the final expression of Eq. (6), integrals with a doubled propagator occur, which in any case have to be evaluated in all cases by IBP identities, or otherwise, independently from the integrand-level reduction discussed in this paper.
3.1.4 Non-planar double box topologies
Figure 4: Feynman graphs contributing to the non-planar double-box numerator under study. There are four contributions, considering gluons and (anti-)ghosts running within the loops.
Here we focus on the numerator of the non-planar double-box topology constructed by the seven Feynman graphs drawn in Fig. 4. The depicted graphs are part of the full-color scattering amplitude of the process $gg\to gg$. The inverse propagators describing the family of this topology are chosen as
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{3}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+k_{2}-p_{3}\right){}^{2},\quad\\ D_{5}=\left(k_{1}+k_{2}\right){}^{2},\quad D_{6}=k_{2}^{2},\quad D_{7}=\left(k_{2}-p_{123}\right){}^{2},\quad D_{8}=\left(k_{2}-p_{12}\right){}^{2},\quad D_{9}=\left(k_{2}-p_{1}\right){}^{2}\end{gathered}$
We seek again to solve Eq. (18) using cut equations and the known analytic expression of the numerator. The top sector cut equations
$D_{1}=D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=0$ (42)
result in the determination of 7 invariants
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\\ k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{2}\to k_{1}\cdot p_{3}-k_{2}\cdot p_{1},\quad k_{2}\cdot p_{3}\to-\left(k_{1}\cdot p_{3}\right)\end{gathered}$ (43)
By applying the above relation on both sides of Eq. (18) we can fully determine the polynomial $P_{7}$. The polynomial $P_{7}$ consists of 97 coefficients over the ISP monomials $\{k_{1}\cdot p_{3},\;k_{1}\cdot\eta,\;k_{2}\cdot p_{1},\;k_{2}\cdot\eta\}$.
Subtracting the $P_{7}$ in Eq. (18), we can now determine the polynomials $P_{6}$ in the same way. There are seven six-cuts and therefore seven $P_{6}$ polynomials to determine. As an example, the first six-cut,
$D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=0$ (44)
leads to
$\begin{gathered}k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\left(k_{2}\cdot p_{1}\right)-k_{2}\cdot p_{2},\quad k_{1}\cdot p_{3}\to k_{2}\cdot p_{1}+k_{2}\cdot p_{2},\\ k_{1}\cdot k_{2}\to k_{1}\cdot p_{1},\quad k_{1}\cdot k_{1}\to-2k_{1}\cdot p_{1},\quad k_{2}\cdot k_{2}\to 0\end{gathered}$ (45)
whereas there are now 5 ISP $\{k_{1}\cdot p_{1},k_{1}\cdot\eta,k_{2}\cdot p_{1},k_{2}\cdot p_{2},k_{2}\cdot\eta\}$. The polynomial $P^{(6)}_{1}$ consists of 126 coefficients. The data for all cuts are summarized in Tab. 1. The analytic solutions for the polynomials satisfy explicitly Eq. (18).
Level Number of cuts Number of coefficients Scaling
7 1 97 5,4,4
6 7 812 4,4,4
5 21 1612 3,3,3
4 35 922 2,2,2
3 35 124 1,1,1
2 21 0 0,0,0
Table 6: Non-planar double box linear fit information beginning with 7 cut.
We now seek to solve Eq. (19), namely projecting over all 9 propagators in the non-planar double-box topology. The top-level cut equations read
$D_{1}=D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=D_{8}=D_{9}=0$ (46)
which leads to
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\\ k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2}\end{gathered}$ (47)
By applying the above relation on both sides of Eq. (19), we can fully determine the polynomial $P_{9}$. The polynomial $P_{9}$ consists of 15 coefficients over the ISP monomials $\{k_{1}\cdot\eta,k_{2}\cdot\eta\}$.
Subtracting as before the $P_{9}$ in Eq. (19), we can now determine the polynomials $P_{8}$ in the same way. There are nine 8-cuts and therefore nine $P_{8}$ polynomials to determine. As an example, the first 8-cut,
$D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D_{7}=D_{8}=D_{9}=0$ (48)
leads to
$\begin{gathered}k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s}{2},\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\\ k_{1}\cdot k_{2}\to k_{1}\cdot p_{1},\quad k_{1}\cdot k_{1}\to-2k_{1}\cdot p_{1},\quad k_{2}\cdot k_{2}\to 0\end{gathered}$ (49)
whereas there are now 3 ISP $\{k_{1}\cdot p_{1},k_{1}\cdot\eta,k_{2}\cdot\eta\}$. The polynomial $P^{(8)}_{1}$ consists of 35 coefficients. The data for all cuts are summarized in Tab. 7. The analytic solutions for the polynomials satisfy explicitly Eq. (19).
Level Number of cuts Number of coefficients Scaling
9 1 15 4,4,4
8 9 267 4,4,4
7 36 1168 3,3,3
6 84 1601 2,2,2
5 126 700 1,1,1
4 126 72 0,0,0
Table 7: Penta-triangle linear fit information beginning with 9-cut.
The rest of the analysis, regarding the fit by cut and the global fit in $d=4-2\epsilon$ and $d=4$ dimensions, is the same as for the planar double box.
3.1.5 $gg\to t\bar{t}$
Figure 5: Feynman graphs contributing to the $gg\to t\bar{t}$ numerator under study. There are three contributions, considering gluons and (anti-)ghosts running within the $k_{1}$ loop, which are collected in the first line of this Figure. In the second line, we quote the tree-order graph used for the summation over polarizations. The thick black line indicates the (anti-)top quark.
In this subsection, we tackle a numerator topology consisting of one massive propagator. More specifically, we study the topology consisting of the three Feynman graphs sketched in Fig. 5, which contribute to the two-loop scattering amplitude of $t\bar{t}$ production. Notice that in order to avoid unnecessary complications due to the appearance of four-dimensional spinors in the helicity amplitude at this stage of the investigation, the analytic expression of the numerator is constructed as fully summed over polarizations of the product of the two-loop contributions with the tree-order one, $\sum({\cal M}^{(0)})^{*}{\cal M}^{(2)}$. The propagators describing the family of this topology are chosen as
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{3}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+k_{2}\right){}^{2},\quad D_{5}=k_{2}^{2},\\ D_{6}=\left(k_{2}-p_{123}\right){}^{2}-m_{t}^{2},\quad D_{7}=\left(k_{2}-p_{12}\right){}^{2},\quad D_{8}=\left(k_{2}-p_{1}\right){}^{2},\quad D_{9}=\left(k_{1}+p_{123}\right){}^{2}\end{gathered}$
The 7-cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot k_{2}\to 0,\\ k_{2}\cdot p_{2}\to\frac{s}{2}-k_{2}\cdot p_{1},\quad k_{2}\cdot p_{3}\to-\frac{s}{2}\end{gathered}$ (50)
and in that case the $P_{7}$ polynomial has no dependence on the loop momenta, $P_{7}=32s^{3}m_{t}^{2}\left(2m_{t}^{2}-s-2t\right)$. The first 6-cut is given by
$\begin{gathered}k_{2}\cdot p_{3}\to-\frac{s}{2},\quad k_{2}\cdot p_{1}\to\frac{s}{2}-k_{2}\cdot p_{2},\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\\ k_{1}\cdot k_{2}\to k_{1}\cdot p_{1},\quad k_{2}\cdot k_{2}\to 0,\quad k_{1}\cdot k_{1}\to-2k_{1}\cdot p_{1}\end{gathered}$ (51)
and the corresponding $P_{6}$ polynomials consists of 15 monomials over $\left\{k_{1}\cdot p_{1},k_{1}\cdot p_{3},k_{2}\cdot p_{2}\right\}$. The data for all cuts are summarized in Tab. 8. The analytic solutions for the polynomials satisfy explicitly Eq. (18).
Level Number of cuts Number of coefficients Scaling
7 1 1 0,0,0
6 7 80 3,3,3
5 21 192 2,2,2
4 35 132 1,1,1
3 35 18 0,0,0
2 21 0 0,0,0
Table 8: $gg\to t\bar{t}$ linear fit information beginning with 7-cut.
As is evident, for processes involving quarks, even massive, the reduction is an order of magnitude simpler than in the case of fully gluonic amplitudes.
Projecting over all 9 propagators in this topology, the 9-cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\\ k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2}\end{gathered}$ (52)
with $P_{9}$ given as before by $P_{9}=32s^{3}m_{t}^{2}\left(2m_{t}^{2}-s-2t\right)$. The data for all cuts are summarized in Tab. 8. The analytic solutions for the polynomials satisfy explicitly Eq. (19).
Level Number of cuts Number of coefficients Scaling
9 1 1 0,0,0
6 9 17 2,1,2
5 36 122 2,2,2
4 84 197 1,1,1
3 126 83 0,0,0
4 126 0 0,0,0
Table 9: $gg\to t\bar{t}$ linear fit information beginning with 9-cut.
In both cases, the rest of the analysis works along the same lines as described for the double-box presented above.
3.2 Five-Point Kinematics
3.2.1 Penta-box topologies
Figure 6: Feynman graphs contributing to the penta-box numerator under study. There are seven contributions, considering gluons and (anti-)ghosts running within the loops.
Here we consider the numerator of the penta-box topology constructed by the seven Feynman graphs depicted in Fig. 6, which contributes to the scattering amplitude of the process $gg\to ggg$. For the analysis presented below, we have used the following helicity assignment: $\lambda_{1}=+1,\lambda_{2}=+1,\lambda_{3}=-1,\lambda_{4}=-1,\lambda_{5}=-1$, for the helicities of the incoming gluons. Our choice for the inverse propagators describing the family of this topology is the following
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{3}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+k_{2}\right){}^{2},\\ D_{5}=k_{2}^{2},\quad D_{6}=\left(k_{2}-p_{1234}\right){}^{2},\quad D_{7}=\left(k_{2}-p_{123}\right){}^{2},\quad D_{8}=\left(k_{2}-p_{12}\right){}^{2},\\ D_{9}=\left(k_{1}+p_{123}\right){}^{2},\quad D_{10}=\left(k_{1}+p_{1234}\right){}^{2},\quad D_{11}=\left(k_{2}-p_{1}\right){}^{2}\end{gathered}$ (53)
The external kinematics is described by 4 independent momenta and five independent invariants, which can be chosen to be $S_{5}=\{s_{12},s_{23},s_{34},s_{45},s_{15}\}$, where $s_{ij}=(p_{i}+p_{j})^{2}$. In order to proceed with the reduction in an analytic setup, the expression of the numerator needs to be expressed in terms of the 11 invariants $V_{11}$, defined as the set of variables $k_{i}\cdot k_{j},$ with $i,j=1,2$, and $k_{i}\cdot p_{j}$ with $i=1,2,\;j=1,\ldots,4$. Since the analytic expression of the numerator involves the polarization vectors of the external gluons, $\varepsilon_{i},i=1,\ldots,5$, scalar products of the form $k_{i}\cdot\varepsilon_{j}$ and $p_{i}\cdot\varepsilon_{j}$ need to be expressed in terms of the set of variables $S_{5}$ and $V_{11}$. To this end, the following relation is used
$q_{i}\cdot q_{j}=G^{-1}_{kl}\,q_{i}\cdot p_{k}\,q_{j}\cdot p_{l}$ (54)
where $q_{i}$ stands for any momentum or polarization vector and $G$ denotes the Gram matrix, $G_{ij}=p_{i}\cdot p_{j},i,j=1,\ldots,4$ expressed in terms of the $S_{5}$ variables. Barring that the analytic expression for the inverse Gram matrix and the numerator are complicated expressions, it is convenient to work in a numerical setup, using exact arithmetic. The numerical values for the variables $S_{5}$, are chosen as $\left\{s_{12}\to 1,s_{34}\to\frac{1}{4},s_{45}\to\frac{1}{4},s_{15}\to-\frac{1}{4},s_{23}\to-\frac{1}{8}\right\}$. The polarization vectors are defined by
$\begin{gathered}\varepsilon_{\mu}^{+}(p_{i})=\frac{1}{\sqrt{2}\bar{u}_{-}(p_{i+1})u_{+}(p_{i})}\bar{u}_{-}(p_{i+1})\gamma_{\mu}u_{-}(p_{i})\\ \varepsilon_{\mu}^{-}(p_{i})=-\frac{1}{\sqrt{2}\bar{u}_{+}(p_{i+1})u_{-}(p_{i})}\bar{u}_{+}(p_{i+1})\gamma_{\mu}u_{+}(p_{i})\end{gathered}$ (55)
for $i=1,\ldots,5$, where in the above formula the following the $p_{6}\to p_{1}$ identification is assumed. With the above identifications, the numerator consists of monomials composed of the $V_{11}$ variables, with exact numerical coefficients.
We seek to solve Eq. (20). The maximal cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s_{12}}{2},\quad k_{2}\cdot k_{2}\to 0,\\ k_{2}\cdot p_{2}\to\frac{s_{12}}{2}-k_{2}\cdot p_{1},\quad k_{2}\cdot p_{3}\to\frac{s_{45}}{2}-\frac{s_{12}}{2},\quad k_{2}\cdot p_{4}\to-\frac{s_{45}}{2}\end{gathered}$ (56)
The $P_{8}$ polynomial consists of 50 terms, composed of monomials in the ISP variables $\left\{k_{1}\cdot p_{3},k_{1}\cdot p_{4},k_{2}\cdot p_{1}\right\}$. By following the usual subtraction procedure, the solution for all polynomials is found to satisfy Eq. (20). The data of this solution are given in Tab. 10.
Level Number of cuts Number of coefficients Scaling
8 1 50 4,5,5
7 8 705 4,4,5
6 28 2550 4,4,4
5 56 3508 3,3,3
4 70 1902 2,2,2
3 56 348 1,1,1
2 28 12 0,0,0
Table 10: Penta-box linear fit information beginning with 8-cut.
Seeking now to solve Eq. (21) the maximal cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s_{12}}{2},\quad k_{1}\cdot p_{3}\to\frac{s_{12}}{2}-\frac{s_{45}}{2},\\ k_{1}\cdot p_{4}\to\frac{s_{45}}{2},\quad k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s_{12}}{2},\\ k_{2}\cdot p_{3}\to\frac{s_{45}}{2}-\frac{s_{12}}{2},\quad k_{2}\cdot p_{4}\to-\frac{s_{45}}{2}\end{gathered}$ (57)
and as before, the cut data are shown in Tab. 11.
Level Number of cuts Number of coefficients Scaling
11 1 1 0,0,0
10 11 47 3,4,4
9 55 502 3,4,4
8 165 2313 3,3,3
7 330 3715 2,2,2
6 462 2255 1,1,1
5 462 425 0,0,0
Table 11: Penta-box linear fit information beginning with 11-cut.
The fit by cut and the global fit in $d=4-2\epsilon$ dimensions is realized as before through the numerical representation of the loop momenta given in section 2.2.
3.2.2 $t\bar{t}H$
Figure 7: Feynman graphs contributing to the $gg\to t\bar{t}H$ numerator under study. There are three contributions, considering gluons and (anti-)ghosts running within the $k_{1}$ loop, which are collected in the first line of this Figure. In the second line, we quote the tree-order graph used for the summation over polarizations. The dashed line indicates the Higgs boson.
Herein, we study a numerator topology consisting of two massive propagators. This corresponds to the three graphs depicted in Fig. 7, which contribute to the scattering amplitude of the process $gg\to ttH$. We again note that in order to avoid unnecessary complications due to the appearance in the helicity amplitude of four-dimensional spinors, the analytic expression of the numerator is constructed as a fully summed over polarizations of the product of the two-loop contributions with the tree-order one. The inverse propagators describing the family for this topology are chosen as
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{3}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+k_{2}\right){}^{2},\quad D_{5}=k_{2}^{2},\\ D_{6}=\left(k_{2}-p_{1234}\right){}^{2}-m_{t}^{2},\quad D_{7}=\left(k_{2}-p_{123}\right){}^{2}-m_{t}^{2},\quad D_{8}=\left(k_{2}-p_{12}\right){}^{2},\\ D_{9}=\left(k_{1}+p_{123}\right){}^{2},\quad D_{10}=\left(k_{1}+p_{1234}\right){}^{2},\quad D_{11}=\left(k_{2}-p_{1}\right){}^{2}\end{gathered}$ (58)
The 8-cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s_{12}}{2},\quad k_{2}\cdot k_{2}\to 0,\\ k_{2}\cdot p_{2}\to\frac{s_{12}}{2}-k_{2}\cdot p_{1},\quad k_{2}\cdot p_{3}\to\frac{1}{2}\left(-m_{t}^{2}-s_{12}+s_{45}\right),\quad k_{2}\cdot p_{4}\to\frac{1}{2}\left(m_{t}^{2}-s_{45}\right)\end{gathered}$ (59)
and in that case, the $P_{8}$ polynomial consists of 32 terms composed of monomials in the ISP variables $k_{1}\cdot p_{3},k_{1}\cdot p_{4},k_{2}\cdot p_{1}$. The data for all cuts are summarized in Tab. 12. The analytic solutions for the polynomials satisfy explicitly Eq. (18).
Level Number of cuts Number of coefficients Scaling
8 1 32 4,3,5
7 8 293 4,3,4
6 28 735 3,2,2
5 56 651 2,2,2
4 70 181 1,1,1
3 56 8 0,0,0
Table 12: $gg\to t\bar{t}H$ linear fit information beginning with 8-cut.
Projecting over all 11 propagators in this topology, the 11-cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-\frac{s_{12}}{2},\quad k_{1}\cdot p_{3}\to\frac{1}{2}\left(s_{12}-s_{45}\right),\\ k_{1}\cdot p_{4}\to\frac{1}{2}\left(s_{45}-m_{t}^{2}\right),\quad k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s_{12}}{2},\\ k_{2}\cdot p_{3}\to\frac{1}{2}\left(-m_{t}^{2}-s_{12}+s_{45}\right),\quad k_{2}\cdot p_{4}\to\frac{1}{2}\left(m_{t}^{2}-s_{45}\right)\end{gathered}$ (60)
with $P_{11}$ consisting of a single term expressed as a function of the external invariants $\left\{m_{H},m_{t},s_{12},s_{15},s_{23},s_{34},s_{45}\right\}$. The data for all cuts are summarized in Tab. 13. The analytic solutions for the polynomials satisfy explicitly Eq. (19).
Level Number of cuts Number of coefficients Scaling
11 1 1 0,0,0
10 11 32 2,2,2
9 55 270 2,2,3
8 165 766 2,2,2
7 330 734 1,1,1
6 462 158 0,0,0
Table 13: $gg\to t\bar{t}H$ linear fit information beginning with 11-cut.
In both cases, the rest of the analysis works in the same lines as for the double-box presented above.
3.3 Six-Point Kinematics
3.3.1 Six-gluon topology
Figure 8: Feynman graph contributing to the six-gluon numerator under study. The normal line is used for denoting a massless quark.
In this subsection, we apply our method to the case of the six-gluon two-loop numerator topology of Fig. 8. This graph is part of the $gg\to gggg$ scattering amplitude. For the analysis presented below, we have used the following helicity assignment: $\lambda_{1}=+1,\lambda_{2}=+1,\lambda_{3}=-1,\lambda_{4}=-1,\lambda_{5}=-1,\lambda_{6}=-1$, for the helicities of the incoming gluons. We define the propagators of the family in which this topology belongs, as
$\begin{gathered}D_{1}=k_{1}^{2},\quad D_{2}=\left(k_{1}+p_{1}\right){}^{2},\quad D_{3}=\left(k_{1}+p_{12}\right){}^{2},\quad D_{4}=\left(k_{1}+p_{123}\right){}^{2},\quad D_{5}=\left(k_{1}+k_{2}\right){}^{2},\\ D_{6}=k_{2}^{2},\quad D_{7}=\left(k_{2}-p_{12345}\right){}^{2},\quad D_{8}=\left(k_{2}-p_{1234}\right){}^{2},\quad D_{9}=\left(k_{2}-p_{123}\right){}^{2},\\ D_{10}=\left(k_{1}+p_{1234}\right){}^{2},\quad D_{11}=\left(k_{1}+p_{12345}\right){}^{2},\quad D_{12}=\left(k_{2}-p_{12}\right){}^{2},\quad D_{13}=\left(k_{2}-p_{1}\right){}^{2}\end{gathered}$
Since out of the six external momenta only four are independent in $d=4$ dimensions, not all 13 propagators in Eq. (3.3.1) are independent. In general, for any $n$-point amplitude, with $n\geq 5$, only 11 propagators are independent. We have chosen the following subset, $\left\{D_{1},\ldots,D_{10},D_{13}\right\}$ having expressed $p_{5}$ in terms of $p_{1},\ldots,p_{4}$ through Eq. (54). As in the case of the penta-box, section 3.2.1, the analytic expressions are hardly manageable, and it is convenient to work in a numerical setup, using exact arithmetic. The numerical values of the invariants are chosen as  [^7] Notice that for 4-dimensional external momenta only 8 invariants are independent. In fact, $s_{345}$, for instance, can be determined by the requirement that the five-momenta Gram determinant, $G(12345)$, vanishes, see also ref. Abreu:2024fei .
$\left\{s_{12}\to 4,s_{23}\to-1,s_{34}\to 1,s_{45}\to\frac{5}{4},s_{56}\to\frac{1}{2},s_{16}\to-1,s_{123}\to 2,s_{234}\to-1,s_{345}\to\frac{31}{12}\right\},$
with $s_{ijk}=(p_{i}+p_{j}+p_{k})^{2}$. For the polarization vectors we follow Eq. (55), with the identification of $p_{7}\to p_{1}$. The 9-cut
$D_{1}=\ldots=D_{9}=0$ (61)
is given, in the numerical point chosen, by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-2,\quad k_{1}\cdot p_{3}\to 1,\quad k_{2}\cdot k_{2}\to 0,\\ k_{2}\cdot p_{2}\to\frac{5}{2}-\frac{5k_{2}\cdot p_{1}}{3},\quad k_{2}\cdot p_{3}\to\frac{2k_{2}\cdot p_{1}}{3}-\frac{3}{2},\quad k_{2}\cdot p_{4}\to-\frac{3}{4}\end{gathered}$ (62)
The reduction data are given in Tab. 14.
Level Number of cuts Number of coefficients Scaling
9 1 21 4,4,6
8 9 355 4,4,6
7 36 1949 4,4,5
6 84 4462 4,4,4
5 126 4540 3,3,3
4 126 2016 2,2,2
3 84 334 1,1,1
2 36 16 0,0,0
Table 14: 6 gluon linear fit information beginning with 9-cut.
Projecting over the set of 11 propagators, referred above, the 11-cut is given by
$\begin{gathered}k_{1}\cdot k_{1}\to 0,\quad k_{1}\cdot k_{2}\to 0,\quad k_{1}\cdot p_{1}\to 0,\quad k_{1}\cdot p_{2}\to-2,\quad k_{1}\cdot p_{3}\to 1,\\ k_{1}\cdot p_{4}\to\frac{3}{4},\quad k_{2}\cdot k_{2}\to 0,\quad k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{5}{2},\\ k_{2}\cdot p_{3}\to-\frac{3}{2},\quad k_{2}\cdot p_{4}\to-\frac{3}{4}\end{gathered}$ (63)
and the reduction data are summarized in Tab. 15.
Level Number of cuts Number of coefficients Scaling
11 1 1 0,0,0
10 11 41 3,3,3
9 55 505 3,3,5
8 165 2365 3,3,4
7 330 4780 3,3,3
6 462 4290 2,2,2
5 462 1592 1,1,1
4 330 200 0,0,0
Table 15: 6 gluon linear fit information beginning with 11-cut.
4 Summary and Discussion
In this paper, we have studied the integrand-level reduction at two loops, solving Eq. (4) and Eq. (17). Let us briefly summarize our findings:
  • We solved the reduction equation by two generic methods, namely, fit by cut and global fit. In the fit by cut approach, we constructed the solutions to the cut equations Eq. (8) and then used them to fit the polynomials $P$ appearing in Eq. (4).
  • A novel element of our studies is the observation that the reduction equation can be extended to include all inverse propagators characterizing a given family and not just the ones appearing in the Feynman graphs under consideration, see Eq. (17). This way, the polynomials depend on a smaller number of variables. Moreover, this approach allows us to use the solutions of the cut equations for all Feynman graphs under the same family.
  • We applied our method in both $d=4-2\epsilon$ and $d=4$ dimensions. In the $d=4-2\epsilon$ case, the numerator is expressed as a function of the four-dimensional external momenta and polarization vectors and the four-dimensional part of the loop momenta along with the $\mu_{ij}$ terms. In all cases, this is readily possible, since the numerator has been known analytically. Cut equations admit a unique solution in terms of the ISP and can be used to fit the coefficients of the polynomials either analytically or numerically.
  • In $d=4$ dimensions, the cut equations do not admit, in general, a unique solution in terms of ISP. We showed, nevertheless, how to construct solutions to both Eq. (4) and Eq. (17).
  • In the global fit approach, both in $d=4-2\epsilon$ and $d=4$ dimensions, solutions to the cut equations are not necessary, but determining the coefficients of the polynomials involves solving linear equations with relatively large matrices. It is though a plausible alternative to the fit by cut approach, especially in validating Eq. (4) and Eq. (17).
  • We have successfully studied the reduction equations for several cases in 4-, 5- and 6-particle topologies, with massless and massive external particles and internal propagators: double-box, penta-triangle, hexa-bubble, non-planar double-box, $t\bar{t}$, penta-box, $t\bar{t}H$ and 6-gluon with a closed quark loop. We conclude that two-loop reduction at the integrand level is within reach for arbitrary processes. We plan to implement the solutions of the cut equations in a numerical setup Ossola:2007ax , in order to allow for a fully-fledged numerical reduction of two-loop integrand numerators.
A crucial element in the direction of the construction of a generic automated two-loop amplitude computational framework based on the integrand-level reduction is the calculation of the dimensionally regulated amplitudes. As we have seen in this paper, the numerical evaluation of the integrands needs to address the dependence on $\mu_{ij}$ and $\epsilon$ terms. This issue is currently under investigation Bevilacqua:2024fec , and we plan to publish our results in the near future new1 .
In order to provide useful results for phenomenological applications, the two-loop Feynman integrals in Eq. (6) need to be provided. At this time, a fully numerical computation of all two-loop integrals Heinrich:2023til ; Liu:2022chg seems quite challenging. Analytic results, on the other hand, are available for master integrals in more and more cases over the recent years Papadopoulos:2015jft ; Gehrmann:2018yef ; Chicherin:2018mue ; Chicherin:2018old ; Abreu:2020jxa ; Canko:2020ylt ; Chicherin:2020oor ; Abreu:2021smk ; Chicherin:2021dyp ; Kardos:2022tpo ; Badger:2022hno ; FebresCordero:2023pww ; Abreu:2023rco ; Abreu:2024yit ; Badger:2024fgb ; Abreu:2024fei ; Henn:2025xrc ; Becchetti:2025qlu ; Bargiela:2025nqc , but IBP solutions in expressing Feynman integrals in terms of master integrals are necessary as well. In principle, this can be achieved in a process-independent way by providing appropriate IBP tables that express the Feynman integrals of Eq. (7) in terms of currently known MIs. We plan to provide a proof of concept of a complete computational framework in the near future new2 , combining HELAC-2LOOP amplitude generator for the numerical evaluation of the numerators in $d=4-2\epsilon$ dimensions, the numerical implementation of the reduction equation, Eq. (17), and the interface to evaluating the currently known Feynman integrals appearing in Eq.( 6), including a detailed analysis of numerical accuracy and performance, paving the way to automated two-loop calculations.
Acknowledgements.
We thank Simon Badger, Hjalte Frellesvig, Adam Kardos, Tiziano Peraro, Vasily Sotnikov, and Yang Zhang for useful discussions. The work of G. Bevilacqua, C. G. Papadopoulos, and A. Spourdalakis is supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the "2nd Call for H.F.R.I. Research Projects to support Faculty Members & Researchers" (Project Number: 02674 HOCTools-II). The work of D. Canko is supported by the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program grant agreement 101040760, High-precision multi-leg Higgs and top physics with finite fields (ERC Starting Grant FFHiggsTop). C. G. Papadopoulos would like to thank the Department of Theoretical Physics at CERN and the Aspen Center for Physics and the Simons Foundation, for their kind hospitality.